# A new indeterminacy-based quantum theory

### v4

@@ 2 Main features: four conditions and two hypotheses
Furthermore, given EnDQT, if a pragmatic reversible decoherence model models specific situations $SI$ with predictive success (like in the case of Wigner’s friend), it allows us to infer the situations $SI$ where a system is interacting with environmental systems that don’t have the DC. Thus, we infer that in situations $SI$ no process that gives rise to determinate values occurs. Thus, note that isolated regions with macroscopic systems inside might isolate these systems from the influence of SDCs. If this is done properly in such a way that we can unitarily manipulate the contents of that region, we have a reversible process inside that region instead of an irreversible one. Then, if some situation, even involving interactions between macroscopic systems (like inside the friend’s lab), is modeled by reversible decoherence with predictive success, we can infer that we have managed to isolate the systems from the influence of SDCs (more on this in the next section). This view is contrary to what is often assumed by relationalist views, such as the MWI, [^27] Wallace(2012). which would consider that determinacy arises within an isolated spatiotemporal region, at least if it is large enough. It is in this sense that, as mentioned in the introduction, for MWI, decoherence is necessary and sufficient for determinacy, and for EnDQT, the interaction with SDCs matters.
@@
Relatedly, a friend could also be shielded from interacting with members of SDCs in the following way. Let’s return to the original example involving systems $A$, $B$, and $C$. As we can see via this example, in order for a system (like $C$) to continue having determinate values of an observable and giving rise to other systems having the DC and having determinate values ("destroying local superpositions"), interactions of the above kind should proceed at other times, i.e., $C$ has to interact with other systems while interacting with $B$. This leads EnDQT to predict a phenomenon that I will call the dissolution of an SDC. If, during the evolution of an SDC, no system interacts with the system like $C$ that is leading the expansion of that SDC, that SDC will disappear, not being able to give rise to further determinate values and the local destruction of superpositions. [^28] Note that system $C$ may also continue having determinate values if its states are decohered by other systems that belong to other SDCs that are expanding. Now, to isolate the friend from the influence of SDCs, we would $just$ need to dissolve the SDCs that could interact with the friend. Note that this is different from isolating systems from interacting with target systems. It’s instead not allowing the target systems to interact with systems with the DC. The phenomenon of dissolution of SDCs is a distinct prediction of EnDQT.
@@
So, what the SDCs-decoherence hypothesis is doing is connecting the pragmatic decoherence models with the fundamental decoherence models, grounding the empirical success of the first on the phenomena represented by the latter. Accepting EnDQT the CDC1)-CDC4), we have good evidence that the SDCs-decoherence hypothesis holds, given the success of models of irreversible decoherence in accounting for measurement-like phenomena, and the success of reversible ones in not accounting for non-measurement-like phenomena. [^29] See, e.g., virtual/reversible decoherence in Schlosshauer(2007). Also, we have seen above how this hypothesis is plausible given that the CDCs lead to phenomena that resemble those represented by pragmatic irreversible decoherence models, which should even resemble them more if the environment is open since the latter will involve even more systems belonging to SDCs.
@@
Contrary to what one may worry about, there is no future dependency or retrocausality for EnDQT in the sense that something giving rise to determinate values in the present depends on how some events turn up to arise in the future, i.e., the interactions turn out to give rise in the future to irreversible decoherence because they become uncontrollable (pragmatically speaking). Instead, given the above SDCs-decoherence hypothesis, our world is such that the pragmatic irreversible decoherence models (in open environments) are a reliable indicator to infer that determinate values arise via local interactions because the environmental systems that participate in this process are such that they have the DC. Also, we have seen how EnDQT justifies the features $i^{*})-iii^{*})$ of fundamental decoherence models, which justify also their success on making inferences about how and when determinate values arise.
@@
Now, it’s time to address the elephant in the room, which is to specify what kind of physical systems initiators are and when we expect the SDCs to start, which corresponds to where initiators began to act. It seems to me that the hypothesis to establish when SDCs started should fulfill the following two desiderata: i’) initiators should have a privileged position that allows them to give rise to widespread SDCs, so that they can explain the standardly assumed widespread existence of systems with determinate values across spacetime, including in the early universe. In other words, it should explain the widespread classicality that we observe or assume to exist. Furthermore, ii’) it should support the SDCs-decoherence hypothesis in the sense that the systems that start SDCs should not be starting SDCs in arbitrary spacetime regions, so that it becomes plausible that sometimes we are able to isolate systems from the influence of SDCs (so that we are able, for example, to recohere the quantum states of a system). It’s also undesirable if we manage to isolate a spacetime region to maintain the quantum systems inside of it in a superposition, and very likely, an initiator could likely manifest itself arbitrarily and destroy those superpositions. So, it’s plausible to consider that initiators should manifest themselves mainly in the early universe since this is the earliest stage when we apply classical physics, in principle it’s a stage that occurred in a clear spacetime region, and it’s plausible to consider that in the early universe, something special happened (more on this below) and that systems in the early universe could be more influential than in the latter stages (more on this below too). So, given i’) and ii’), it’s plausible to assume the following general hypothesis,
@@
At least most current SDCs in our universe started in the early universe, and initiators had a privileged role in this stage in terms of interacting with other quantum systems, which led to the formation of these SDCs(SDCs-starting hypothesis).
@@
Inflation is typically considered to have been driven by a scalar field, called the inflaton. [^30] Some models postulate multiple inflaton fields; for simplicity, here I will just consider that there is one field. It’s hypothesized that the zero-point fluctuations of this quantized scalar field and the associated energy-momentum fluctuations and gravitational field that it gave rise to, when amplified by the rapid expansion of inflation, attracted matter. So, it’s standardly hypothesized that these fluctuations served as a source that gave rise to the cosmic structure in our universe (e.g., galaxy, galaxy clusters, etc.). [^31] Liddle & Lyth (2009). The inflaton field is often treated classically, and the effects of these fluctuations are observed via slight differences in temperature in the Cosmic Microwave Background. The problem to solve is to explain how these quantum fluctuations became classical during the early stages of the evolution of the universe. To my knowledge, the earliest reasonably accepted decoherence processes involved the decoherence of these fluctuations and aimed to address this problem. Many decoherence models were formulated to describe this process. [^32] See, e.g., Boyanovsky (2015); Burgess, Holman, & Tasinato (2015); Burgess, Holman, & Tolley (2015); Burgess, Holman, Kaplanek; Kiefer, Lohmar, Polarski, & Starobinsky (2007); Kiefer & Polarski (2009); Liu, Sou, & Wang (2016); Martin & Vennin (2016, 2020, 2022); Martin, Vennin, & Starobinsky (2021); and Nath Raveendran, Parattu, & Sriramkumar (2022). In a toy model, $\left|s_{i}\right\rangle_{S}$ in eq.(2) could be the field amplitude/fock state with momentum k of the inflaton field fluctuations $S$. $E$ would be the environment (more on this below). [^33] See, e.g., Kiefer & Polarski (2009) and references therein for more details on these decoherence models. We could invoke decoherence plus the MWI to perhaps make these models more satisfactory. Alternative proposals appealed to spontaneous collapse theories to explain how quantum fluctuations become classical. [^34] See, e.g., Perez et al. (2006). However, we can instead appeal to the more conservative approach proposed here.
@@
Given the above heuristic, we can consider that the systems implicated in this process are the initiators, having a series of attractive features that fulfill the desiderata aimed by the SDCs-starting hypothesis. First, it fulfills the desiderata i) because of its influential position and role mentioned above that gave rise to the cosmic structure. Furthermore, the decay of the inflaton in the reheating stage is often hypothesized to have given rise at least to ordinary matter. Second, it fulfills the desiderata ii) given that the inflaton field doesn’t seem to manifest itself in the present universe, we can, in principle, build arguments that consider that this field in the later stages of the universe didn’t give rise anymore to SDCs or negligibly so. For instance, in the so-called reheating phase, it’s standardly considered that the inflaton field reached the absolute minimum of its potential and stayed there (and has been staying there). If such minimum corresponds to the point where the field is zero or approximately zero (see Martin et al. (2014) for some empirically supported potentials by the Planck satellite that fulfill this requirement), and if we consider that the coupling of the inflaton field to all other fields/systems in the Lagrangian density that describes or governs our universe depends on the value of the inflaton field in such a way that the interaction terms are zero (or approximately so) when the field zero, we can consider that the inflaton field in the stages of the evolution of the universe after the reheating phase (which includes the phase where we are now) will at least rarely interact with other fields/systems. So, it will (at least) rarely give rise to SDCs in these later stages. [^35] One may worry that in other stages of the evolution of the universe (to put it in very rough and intuitive terms), there may be the creation of virtual particles-antiparticles pairs from the vacuum occupied by the inflaton field. These particles may give rise to SDCs. This isn’t, in principle, a problem. In many inflaton models (see, e.g., Binétruy & Dvali, 1996; Halyo, 1996; Lyth & Riotto, 1999; McDonald, 2000), the inflaton is considered to be very massive/energetic, and so those particles will be too short-lived (see Roberts & Butterfield (2020) for a rigorous explanation for why this is the case) to be able to give rise to SDCs at least significantly. Note that the particles that arise from the vacuum also have to be able to decohere other quantum systems to give rise to SDCs.
@@
Let’s represent the Lagrangian density of our universe obeying the above constraints as $\mathcal{L}_{SDC}$. Since it fulfills the desiderata i) and ii), I will call this hypothesis that appeals to inflation to explain the beginning of SDCs and the inflaton as the initiators, the inflationary-starting hypothesis, and could be stated in the following way,
@@
Which systems are the initiators, more concretely? For example, Kiefer & Polarski (2009) list various possible environmental systems such as some other quantum fields [^36] Possibly coming from string theory. or $parts$ of the fluctuations themselves. [^37] E.g., an environment involving modes of the field with different momentum k. If we adopt fluctuations themselves as initiators, they will have the DC concerning any system and give rise to other systems with determinate values by interacting with other fluctuations and so on, forming SDCs. Note that if we instead consider that other fields have the DC, being these fields the initiators, we would need to explain why these other fields aren’t currently still giving rise to SDCs in agreement with the SDCs-starting hypothesis.
Instead of adopting the simple initiators that I have been talking about, the second possibility involves postulating another kind of initiator, which I will call reactive initiators. Contrary to non-initiator systems,
Reactive initiators S are systems that, upon specific interactions with other systems E that don’t have the DC, acquire the DC concerning any system due to E and a determinate value during interactions, where these interactions can be inferred via the pragmatic irreversible decoherence of S by E.
@@
The inflaton field, with its fluctuations, would be a reactive initiator. The interactions with these perturbations represented via decoherence, where the environmental systems are other fields, give rise to these systems having the DC and a determinate value.
@@
As we can see, the SDCs-starting hypothesis, offers us a more precise means of establishing initiators’ locations. With the inflationary-starting hypothesis, I have pointed out that we already have models that agree with the above hypotheses and simply appeal to the dominant paradigm in cosmology.
@@
Furthermore, appealing to the inflaton as the initiator provides various advantages to EnDQT and possible scientific and philosophical payoffs. First, it shows that EnDQT is a more parsimonious and perhaps more explanatory theory than other interpretations of quantum theory. On top of their ontological or mathematical additions to physics and quantum theory, other quantum theories very likely need to postulate the inflaton field as belonging to the initial conditions of the universe and interpret it to address the problems that inflation is meant to address or as an instance of the so-called past hypothesis to explain the temporal asymmetries. [^38] Albert (2000). This hypothesis puts a constraint on the state of the early universe so that the dynamics of the physical states in the future have the required temporal asymmetries, postulating a time-asymmetric boundary condition. Following Wallace (2023), he considered that "[t]he Past-Hypothesis, (…) is that the world came into being (or at least coalesced out of Planck-scale physics) in the local quantum vacuum state for a homogeneous, isotropic, inflationary spacetime." This vacuum state is also called the Bunch-Davies vacuum, [^39] Bunch & Davies (1978). which is considered to be the initial state of the fluctuations of the inflaton field. [^40] More concretely, it is the minimum energy eigenstate of the Hamiltonian for the primordial fluctuations infinitely back in the past. However, EnDQT doesn’t need those additional mathematical or ontological postulates. They are already a part of how EnDQT regards the inflaton as an initiator and the interactions that arise from there, which can be regarded as an integral part of the theory. [^41] Even if one rejects inflation or the initial state of the inflaton as the past hypothesis, one needs to provide a substitute to these hypotheses that provide the same explanations, which is a tall order.
@@
Also, EnDQT offers the possibility of explaining or grounding the features of the initial conditions and the phenomena that they explain in something more fundamental. Quantum phenomena itself is arguably more fundamental than phenomena described by classical cosmology, particle physics, thermodynamics, and statistical mechanics. The consequences of inflation, which involves the homogeneity of (relativistically speaking) causally separated different regions of space, perhaps the different temporal asymmetries, the geometry of the universe to be nearly flat, etc., would be seen as a consequence of or as emergent phenomena coming from the initiators that allow for SDCs to spread across the universe and that don’t manifest themselves anymore. If one doesn’t regard the initiators and their consequences as explaining these higher-level phenomena, at least EnDQT can be seen as reducing the mystery of these initial conditions to a more fundamental mystery.
@@
Second and relatedly, EnDQT reduces the at-first-sight ad hocness of the inflaton field by providing a more fundamental role for it, which is solving problems with the more fundamental theory, quantum theory, rather than just solving some less fundamental problems.
@@
Third, the inflationary-SDCs hypothesis offers predictions, which might allow us to constrain the very permissive models of inflation further. [^42] Ijjas et al., (2013) and Dawid & McCoy (2023). For instance, as was mentioned, one of the constraints on inflationary models, which can be regarded as a prediction of EnDQT, is that the inflationary potential should have an absolute minimum when its value is zero in such a way that the Lagrangian turns off the interactions of the inflaton field with other fields/systems once the inflationary potential reaches its absolute minimum.
@@
Fourth, as I have mentioned, instead of explaining how primordial fluctuations arose from quantum fluctuations of the inflaton field by appealing to spontaneous collapse theories (Perez et al., 2006) that modify the equations of QT, or MWI and decoherence, which leads to the problem of quantum probabilities, EnDQT provides a conservative solution that doesn’t suffer from the above potential issues.
@@
One might object that EnDQT, with its initiators, is built on top of speculative hypotheses. However, note that the above hypotheses are as speculative as any other hypotheses postulated by the currently popular interpretations of QT, and they turn out to be conservative in the sense that they don’t involve any modification of standard QT. [^43] MWI supporters might claim they make no speculative hypotheses, but a realist attitude to whatever lies beneath the multiplicity of worlds is itself a speculative hypothesis. Furthermore, given the benefits of achieving UT), LC), and NR) (as we will see), it’s worth taking the above hypotheses of EnDQT seriously. Also, I have shown that the inflationary-starting hypothesis provides a series of philosophical and scientific payoffs. So, overall, EnDQT yields worthy payoffs. On top of that, the beginning of SDCs occurs in regions of spacetime where we expect special events to occur. Finally, I should emphasize that every approach to QT so far needs to appeal to special initial conditions for one reason or another. However, EnDQT does that while providing additional benefits.

### v5

@@ 2.2 Two hypotheses
So, note that isolated regions with macroscopic systems inside might isolate a macroscopic system $S$ from the influence of SDCs (like in the Wigner’s friend situation). Suppose this is done properly so that we can unitarily manipulate the contents of that region. In that case, we have a reversible process inside that region instead of an irreversible one. Thus, if we have some situation that even involves interactions between macroscopic systems (like inside the friend’s lab) but that is modeled by reversible decoherence with predictive success, we can infer that we have managed to isolate the systems from the influence of SDCs (more on this in Section 3). Whether we are able to achieve this isolation depends fundamentally on the history of the SDCs (modeled via fundamental decoherence models) that could interact with $S$ in those situations and on our ingenuity in shielding it from interacting with SDCs (more on this below).
This view held by EnDQT contrasts with the one often assumed by relationalist views, such as the MWI,434343[95]. which would consider that determinacy arises within a large enough isolated spatiotemporal region with systems decohering each other inside of it. It is in this sense that, as mentioned in the introduction, for MWI, decoherence is necessary and sufficient for determinacy, and for EnDQT, the interaction with SDCs matters.
@@
Relatedly, a friend could also be shielded from interacting with members of SDCs in the following way. Let’s return to the original example involving systems $A$, $B$, and $C$. As we can see via this example, in order for a system (like $C$) to continue having determinate values of an observable and giving rise to other systems having the DC and having determinate values, interactions of the above kind should proceed at other times, i.e., $C$ has to interact with other systems while interacting with $B$. This leads EnDQT to predict a phenomenon that I will call the dissolution of an SDC. If, during the evolution of an SDC, no system interacts with the system like $C$ that is leading the expansion of that SDC, that SDC will disappear, not being able to give rise to further determinate values and the local destruction of superpositions. [^44] Note that system $C$ may also continue having determinate values if its states are decohered by other systems that belong to other SDCs that are expanding. Now, to isolate the friend from the influence of SDCs, we would $just$ need to dissolve the SDCs that could interact with the friend. Note that this is different from isolating systems from interacting with target systems. It is instead not allowing the target systems to interact with systems with the DC. The phenomenon of dissolution of SDCs is a distinct prediction of EnDQT.
@@
So, what the SDCs-decoherence hypothesis is doing is connecting the pragmatic decoherence models with the fundamental decoherence models, grounding the empirical success of the first on the phenomena represented by the latter. Accepting EnDQT and the CDC1)-CDC4), we have good evidence that the SDCs-decoherence hypothesis holds, given the success of models of irreversible decoherence in accounting for measurement-like phenomena, and the success of reversible ones in not accounting for non-measurement-like phenomena. [^45] See, e.g., virtual/reversible decoherence in [89]. Also, we have seen above how this hypothesis is plausible given that the CDCs lead to phenomena that resemble those represented by pragmatic irreversible decoherence models.
@@
Contrary to what one may worry about, there is no future dependency or retrocausality for EnDQT in the sense that something giving rise to determinate values in the present depends on how some events turn up to arise in the future, i.e., the interactions turn out to give rise in the future to irreversible decoherence because they become uncontrollable (pragmatically speaking). Instead, given the above SDCs-decoherence hypothesis, it is hypothesized that our world is such that the empirically successful local pragmatic irreversible decoherence models (in open environments) are a reliable indicator to infer that determinate values arise via local interactions because the environmental systems that participate in this process are such that they have the DC.
@@
Now, it is time to address the elephant in the room, which is to specify what kind of physical systems initiators are and when we expect the SDCs to start, i.e., when initiators begun to act. It seems to me that the hypothesis to establish when SDCs started should fulfill the following two desiderata: i’) initiators should have a privileged position that allows them to give rise to widespread SDCs so that they can explain the standardly assumed widespread existence of systems with determinate values across spacetime, including in the early universe. In other words, it should explain the widespread classicality that we observe or assume to exist. Furthermore, ii’) it should support the SDCs-decoherence hypothesis in the sense that the systems that start SDCs should not be starting SDCs in arbitrary spacetime regions so that it becomes plausible that sometimes we are able to isolate systems from the influence of SDCs (so that we are able, for example, to recohere the quantum states of a system). It is also undesirable that, if we manage to isolate a spacetime region to maintain the quantum systems inside it in a superposition, an initiator could likely manifest arbitrarily and destroy those superpositions. This desideratum is particularly important in the case of initiators of the kind a). These initiators always have the DC concerning any system.
It is plausible that initiators should manifest themselves mainly in the early universe since this is the earliest stage when we apply classical physics. Furthermore, it is plausible to consider that in the early universe, something special happened (more on this below) and that systems in the early universe could be more influential than in the latter stages (more on this below, too). So, given i’) and ii’), it is plausible to assume the following general empirical hypothesis as one possibility for when SDCs started,
@@
At least most current SDCs in our universe started in what we consider the early universe, and initiators had a privileged role in this stage in terms of interacting with other quantum systems, which led to the formation of these SDCs (SDCs-starting hypothesis).
@@
Inflation is typically considered to have been driven by a scalar field $\phi$, called the inflaton. [^46] Some models postulate multiple inflaton fields; for simplicity, here I will just consider that there is one field. It is hypothesized that the zero-point fluctuations of the quantized inflaton scalar field in some regions (i.e., fluctuations of the field in the vacuum state) and the associated energy-momentum fluctuations and gravitational field, amplified by the rapid expansion of inflation, attracted more matter than in other regions. Then, it is hypothesized that this phenomenon gave rise to the (non-evenly distributed) cosmic structure in our universe (e.g., galaxy, galaxy clusters, etc.).474747[60]. The inflaton field is often treated classically, and the effects of these fluctuations are observed via slight temperature anisotropies in the Cosmic Microwave Background. The problem is to explain how these quantum fluctuations became classical during the early stages of the evolution of the universe. To my knowledge, the earliest reasonably accepted decoherence processes involved the decoherence of these fluctuations and aimed to address this problem. Many decoherence models were formulated to describe this process. [^48] See, e.g., [15, 21, 19, 57, 62, 20, 85].
In this article, I am not going to enter into details concerning how decoherence helps address this problem; however, we can briefly see how that can be done. In a toy model, each quantum state $\left|s_{i}\right\rangle_{S}$ in eq.(5) could be the field amplitude/fock state with momentum k of the inflaton field fluctuations $S$ in a spacetime region, and thus $i$ would be the particle occupation number (i.e., the number of particles occupying a given mode k ). $E$ would be the environment (more on this below). [^49] See, e.g., [56] and references therein for more details on these decoherence models. We could invoke decoherence plus the MWI to perhaps make these models more satisfactory. Alternative proposals appealed to spontaneous collapse theories to explain how quantum fluctuations become classical. [^50] See, e.g., [80]. However, we can instead appeal to the more conservative approach proposed here.
@@
Given the above heuristic, we can assume that currently the inflaton is a plausible possible candidate to be an initiator. This assumption has a series of attractive features that fulfill the desiderata aimed by the SDCs-starting hypothesis. First, it fulfills the desiderata i’) because of its influential position and role mentioned above, i.e., it gives rise to the cosmic structure. Furthermore, the decay of the inflaton in the so-called reheating stage is often hypothesized to have given rise to, at least, ordinary matter. Second, it fulfills the desiderata ii’) given that the inflaton field doesn’t seem to manifest itself in our present universe, we can, in principle, build arguments that consider that this field in the later stages of the universe didn’t give rise anymore to SDCs or negligibly so. For instance, in the so-called reheating phase, it is standardly considered that the inflaton field reached the absolute minimum of its potential $V(\phi)$ and stayed there (and has been staying there). [^51] At least in our universe. Let’s assume that such minimum corresponds to the point where the field is zero or approximately zero (see [63] for some empirically supported potentials by the Planck satellite that fulfill this requirement). Let’s also consider that the coupling of the inflaton field to all other fields/systems in the Lagrangian density that describes or governs our universe depends on the value of the inflaton field in such a way that the interaction terms are zero (or approximately so) when the field zero. Given these two assumptions, we can consider that the inflaton field in the stages of the evolution of the universe after the reheating phase (which includes the phase where we are now) will at least rarely interact with other fields/systems. So, if we assume initiators of the kind a), it will (at least) rarely give rise to SDCs in these later stages. [^52] One may worry that in other stages of the evolution of the universe (to put it in very rough and intuitive terms), there may be the creation of virtual particles-antiparticles pairs from the vacuum occupied by the inflaton field. These particles may give rise to SDCs. This isn’t, in principle, a problem. In many inflaton models (see, e.g., [13, 46, 61, 67], the inflaton is considered to be very massive/energetic, and so those particles will be too short-lived (see [87] for a rigorous explanation for why this is the case) to be able to give rise to SDCs at least significantly. Note that the particles that arise from the vacuum also have to be able to decohere other quantum systems to give rise to SDCs.
@@
Let’s represent the Lagrangian density of our universe obeying the above desiderata, and whose initiator is the inflaton field, as $\mathcal{L}_{SDC}$. Since it fulfills the desiderata i’) and ii’), I will call this hypothesis that appeals to inflation to explain the beginning of SDCs and the inflaton as the initiator, the inflationary-starting hypothesis. This hypothesis could be stated in the following way,
@@
How can we understand what are these initiators more concretely? This is largely an open question, and there are multiple possibilities. It is important that these possibilities concern local interactions. For example, [56] list various possible environmental systems that decohere the fluctuations of the inflaton field, such as some other quantum fields [^53] Possibly coming from string theory. or $parts$ of the fluctuations themselves. [^54] E.g., an environment involving modes of the field with different momentum k. If we adopt the inflaton and its fluctuations as initiators, they will have the DC concerning any system and give rise to other systems with determinate values by interacting with (i.e., decohering) other fluctuations and so on, forming SDCs. Note that if we consider instead that other fields have the DC, being these fields the initiators, we would need to explain why these other fields aren’t currently still giving rise to SDCs in agreement with the SDCs-starting hypothesis.
The description of the inflaton field and the rest of the fields interacting with it needs a quantum field theoretical treatment. In this article, I haven’t shown how EnDQT can be understood in the context of quantum field theory, but in principle, it won’t be problematic to provide such treatment.555555[82]. Briefly, in this case, since we are interested in local interactions, we only consider quantum fields in single localized bounded spacetime regions. These are the fundamental quantum systems that transmit the DC. Briefly, in one possible approach to understanding these systems, we associate to a quantum field $\phi(x,t)$ in a finite spacetime region, such as the inflaton field in the spatial region $x$ at $t$, a wavefunctional $\Psi[\phi,t]$. The latter assigns a complex amplitude to each possible configuration of classical fields in spacetime regions, leading to a superposition of these configurations. A quantum field, which concerns a localized finite region, and the determinate values that may arise in that region will also be represented by the observables that act on the wavefunctional in that region. So, the DC is transmitted between quantum fields occupying a single localized region via local interactions. They are local because the observables that represent these systems concern local spacetime regions and interactions. [^56] We guarantee this via conditions such as the so-called microcausality conditions, clustering decomposition ([99]), and the Hamiltonian being local. Thus, it’s expected that systems have determinate values and spread the DC in local regions of spacetime. The inflaton field is the initiator, but it transmits the DC via local interactions.
How can we understand what are these initiators more concretely? This is largely an open question, and there are multiple possibilities. It is important that these possibilities concern local interactions. For example, [56] list various possible environmental systems that decohere the fluctuations of the inflaton field, such as some other quantum fields [^57] Possibly coming from string theory. or $parts$ of the fluctuations themselves. [^58] E.g., an environment involving modes of the same field as the target system, but with different momentum k. If we adopt the inflaton (with its fluctuations) as initiators, they give rise to other systems with determinate values by interacting with them and so on, forming SDCs. [^59] Note that if we adopt initiators of the kind a) and consider instead that other fields have the DC, being these fields the initiators, we would need to explain why these other fields aren’t currently still giving rise to SDCs in agreement with the SDCs-starting hypothesis. Instead of adopting the simple initiators I have been talking about, an alternative possibility involves postulating another kind of initiator, which I will call reactive initiators. Contrary to non-initiator systems, reactive initiators S are systems that, upon specific interactions with other systems E that don’t have the DC, acquire the DC either concerning an appropriate system in agreement with CDC1) or any system, and a determinate value during interactions. These interactions can be inferred via the pragmatic irreversible decoherence of $S$ by $E$.
@@
In this case, the inflaton field system, with its fluctuations, would be a reactive initiator. The interactions with these perturbations, where the environmental systems would be other fields, would give rise to the fluctuations in a region having the DC.
@@
As we can see, the SDCs-starting hypothesis offers resources to establish what initiators may be. With the inflationary-starting hypothesis, I have pointed out that we already have models that agree with the above hypotheses and that simply appeal to the dominant paradigm in cosmology.
@@
Furthermore, appealing to the inflaton as the initiator provides various advantages to EnDQT and possible scientific and philosophical payoffs. First, it potentially shows that EnDQT is a more parsimonious theory than other QTs. [^60] Of course, it becomes less parsimonious if we adopt the initiator of the kind a). On top of their ontological or mathematical additions to physics and quantum theory, other QTs very likely need to postulate the inflaton field (or some analog to this field) as belonging to the initial conditions of the universe and interpret it to address the problems that inflation is meant to address or as an instance of the so-called past hypothesis to explain the temporal asymmetries.616161[1]. This hypothesis constrains the state of the early universe so that the dynamics of the physical states in the future have the required temporal asymmetries, postulating a time-asymmetric boundary condition. Following [97], we consider that “[t]he Past-Hypothesis, (…) is that the world came into being (or at least coalesced out of Planck-scale physics) in the local quantum vacuum state for a homogeneous, isotropic, inflationary spacetime.” This vacuum state is also called the Bunch-Davies vacuum,626262[18]. which is considered to be the initial state of the fluctuations of the inflaton field. [^63] More concretely, it is the minimum energy eigenstate of the Hamiltonian for the primordial fluctuations infinitely back in the past. However, EnDQT doesn’t need those additional mathematical or ontological postulates. They are already a part of how EnDQT regards the inflaton as an initiator and the interactions that arise from there, having a fundamental role in the theory. [^64] Even if one rejects inflation or the initial state of the inflaton as the past hypothesis, one needs to provide a substitute to these hypotheses that provide the same explanations, which is a tall order.
@@
Second, it may a) show that EnDQT is also a more explanatory theory than other interpretations of quantum theory, or b) it at least offers a better scientific reductionist approach toward the above physical phenomena, which is often seen as valuable. More concretely, given the fundamental role of the inflaton mentioned above, EnDQT offers the possibility of explaining, or at least reducing the features of the initial conditions and the phenomena they aim to explain, together with the laws of physics, in something more fundamental. Buying the EnDQT picture, the story would go roughly like this: fundamental features of quantum phenomena involving initiators and their behavior are arguably more fundamental than phenomena described by classical cosmology, particle physics, thermodynamics, and statistical mechanics. The consequences of inflation (or any physical entity with a similar role), which is believed to involve the homogeneity of (relativistically speaking) causally separated different regions of space, perhaps the different temporal asymmetries, the geometry of the universe to be nearly flat, the seeds of structure formation, etc., and perhaps the state given by the Past Hypotheses, would be seen as fundamentally arising from an initiator with certain features plus the laws of physics that describe/govern its behavior and interactions. These features allowed for SDCs to spread throughout the universe in a specific way but also led the initiator to not manifest itself anymore, not currently giving rise to SDCs (being in the absolute minimum of its potential).
Note that there is a sense in which this could be an explanation for the initial conditions of the universe because non-fundamental $special$ facts about the initial conditions of the universe can be grounded on the fundamental $special$ facts about QT. This is because initiators as special entities, and the phenomena that they give rise to, are fundamental for EnDQT and QT. There is much more to say about this. Whether one should regard the initiators and their features as explaining some features of the initial conditions and their consequences, or at least successfully reducing this to a more fundamental mystery, is a topic for future work. However, for now, note that it is plausible to consider that EnDQT offers this interesting possibility.
So, even though EnDQT possibly ties a solution to the measurement problem in some features in the early universe, and that might be considered a problematic approach, any interpretation still has to postulate such initial conditions or events in the early universe. Furthermore, EnDQT can allow for other important advantages when it ties these initial conditions to initiators, which is a benefit of this approach.
@@
Third, EnDQT diminishes the at-first-sight ad hocness of the inflaton field by providing it with a more fundamental role. This role involves solving problems with the more fundamental theory, quantum theory, rather than just solving some less fundamental problems.
@@
Fourth, the inflationary-SDCs hypothesis offers predictions, which might allow us to constrain the very permissive models of inflation further.656565[51, 30]. For instance, as was mentioned, one of the possible constraints on inflationary models that EnDQT imposes is that the inflationary potential should have an absolute minimum when its value is zero in such a way that the Lagrangian turns off the interactions of the inflaton field with other fields/systems once the inflationary potential reaches its absolute minimum. This can be regarded as a prediction of EnDQT.
@@
Fifth, as I have mentioned, instead of explaining how primordial fluctuations arose from quantum fluctuations of the inflaton field by appealing to spontaneous collapse theories [80] that modify the equations of QT, or MWI and decoherence, which leads to the problem of quantum probabilities, EnDQT provides a conservative solution that doesn’t suffer from the above potential issues.
@@
One might object that I didn’t provide details about how these inflation models work to form the SDCs, and so the above hypothesis is not well-supported. This is very much a topic for future work. However, notice that as long as one believes in inflation, this shouldn’t be a worry. This is because as long as inflationary models are well-developed and allow for local interactions, working out these models shouldn’t be an issue (I have mentioned some decoherence models above that may be adapted to an SDC scenario like in the Appendix A). If it is an issue, it is also an issue to the inflationary paradigm and not just a problem for EnDQT. One might object that EnDQT, with its initiators, is built on top of speculative hypotheses. However, note that the above hypotheses are as speculative as any other hypotheses postulated by the currently popular interpretations of QT, and they turn out to be conservative in the sense that they don’t involve any modification of standard QT. [^66] MWI supporters might claim they make no speculative hypotheses, but a realist attitude to whatever lies beneath the multiplicity of worlds is itself a speculative hypothesis. Furthermore, given the benefits of achieving UT), LC), and NR) (as we will see), it is worth taking the above hypotheses of EnDQT seriously. Also, I have shown that the inflationary-starting hypothesis provides a series of philosophical and scientific payoffs. So, overall, EnDQT yields worthy payoffs. On top of that, the beginning of SDCs occurs in regions of spacetime where we expect special events to occur. Finally, I should emphasize that every approach to QT so far needs to appeal to special initial conditions for one reason or another. However, EnDQT does that while providing additional benefits.